How many electrons are transferred in each of these 2 reactions?
1. hcooh--> co2+h2
2. hcooh+h2o-->ch3oh+o2
Answer #1

Solution:

1) HCOOH = CO2 + H2

The oxidation number of carbon in HCOOH is +2 and in CO2, it is +4.

HCOOH (+2) ===== CO2 (+4)

Hence, number of electron transfered = 4 - 2 = 2

Therefore, process is oxidation.

2) HCOOH + H2O ===== CH3OH + O2

The oxidation number of C in HCOOH is +2 and -2 in CH3OH

Therefore,

HCOOH ( +2) ===== CH3OH (-2)

Thus,

Number of electron transfered = -2 - (+2) = -4

Hence, 4 electrons are transfered and procwss in reduction.